# Using negative controls to identify causal effects with invalid instrumental variables

### v4

@@ A.10 The average treatment effect in the treated
$\beta(\mathbf{C})$ $=\frac{\delta^{Y}(\mathbf{C})-t(1,\mathbf{C})}{\delta^{A}(\mathbf{C})}$
$\delta^{Y}(\mathbf{C})$ $=E(Y|Z=1,S=1,\mathbf{C})-E(Y|Z=0,S=1,\mathbf{C})$
$\delta^{A}(\mathbf{C})$ $=f(A=1|Z=1,S=1,\mathbf{C})-f(A=1|Z=0,S=1,\mathbf{C}).$
@@ Appendix B Proofs
$E(Y|A=a,z,S=1,\mathbf{c})=$ $\beta(a,z,\mathbf{c})+\gamma(z,\mathbf{c})\{a-f(A=1|z,S=1,\mathbf{c})\}$
@@
$t(z,\mathbf{c})$ $=E(Y^{0}|Z=1,S=0,\mathbf{C}=\mathbf{c})-E(Y^{0}|Z=0,S=0,\mathbf{C}=\mathbf{c})$ (Assumption 4)
@@
$=\beta(a,z,\mathbf{c})+\gamma(z,\mathbf{c})\{a-f(A=1|z,S=1,\mathbf{c})\}$
@@
$\epsilon_{1}$ $=Y-\beta(\mathbf{A,Z,C})-[q(\mathbf{A,Z,C})-E\{q(\mathbf{A,Z,C})|\mathbf{Z},S=1,\mathbf{C}\}]-b^{*}_{1}(\mathbf{C})-t(\mathbf{Z,C})$
$\epsilon^{*}_{1}$ $=\epsilon_{1}+[q(\mathbf{A,Z,C})-E\{q(\mathbf{A,Z,C})|\mathbf{Z},S=1,\mathbf{C}\}]$
$\epsilon_{0}$ $=Y-b_{0}(\mathbf{C})-t(\mathbf{Z,C})$
@@
$c_{2}(\mathbf{A,Z},\mathbf{C})=[\phi_{1}(\mathbf{Z,C})-E\{\phi_{1}(\mathbf{Z,C})|S=1,\mathbf{C}\}][q(\mathbf{A,Z,C})-E\{q(\mathbf{A,Z,C})|\mathbf{Z},S=1,\mathbf{C}\}]$
@@
=$=\bigg{\{}S\epsilon^{*}_{1}[\phi_{1}(\mathbf{Z,C})-E\{\phi_{1}(\mathbf{Z,C})|S=1,\mathbf{C}\}]:\textit{$\phi_{1}(\mathbf{Z,C})$ unrestricted}\bigg{\}}\cap L_{2}^{0}$
@@
$\Lambda^{\perp}_{nuis_{2}}\cap\Lambda^{\perp}_{nuis_{8}}=\left\{(1-S)\epsilon_{0}[\phi_{2}(\mathbf{Z,C})-E\{\phi_{2}(\mathbf{Z,C})|\mathbf{C},S=0\}];\textit{$\phi_{2}(\mathbf{Z,C})$ unrestricted}\right\}$
@@
$\quad+(1-S)\epsilon_{0}[\phi_{2}(\mathbf{Z,C})-E\{\phi_{2}(\mathbf{Z,C})|\mathbf{C},S=0\}]:\textit{$\phi_{1}(\mathbf{Z,C}),\phi_{2}(\mathbf{Z,C})$ unrestricted}\bigg{\}}.$
@@
$=E\left[\phi(\mathbf{Z},S,\mathbf{C};\tau^{\dagger},\alpha^{\dagger},\rho^{\dagger})\{b_{0}(\mathbf{C})+b_{1}(\mathbf{C})S+t(\mathbf{Z,C})-b_{0}(\mathbf{C};\theta_{0}^{\dagger})-b_{1}(\mathbf{C};\theta_{1}^{\dagger})S-t(\mathbf{Z,C};\nu^{\dagger})\}\right]=0$
@@
where the final equality follows since $f(\mathbf{Z}|S,\mathbf{C})=f(\mathbf{Z}|S,\mathbf{C};\nu^{\dagger},\rho^{\dagger})$.
@@
where the final equality follows since $f(S|S,\mathbf{Z,C})=f(S|\mathbf{Z,C};\alpha^{\dagger},\rho^{\dagger})$.
@@
$E(Y|A=a,z,S=1,\mathbf{c})$
$=E(Y^{1}-Y^{0}|A=1,z,S=1,\mathbf{c})a$
$\quad+\gamma(z,\mathbf{c})\{a-Pr(A=1|z,S=1,\mathbf{c})\}$
@@
$E\{Y-E(Y|Z=z,S=0,\mathbf{c})+E(Y|Z=0,S=0,\mathbf{c})|z,S=1,\mathbf{c}\}$
@@
$\Lambda^{\perp}_{nuis_{2}}\cap\Lambda^{\perp}_{nuis_{8}}=\left\{(1-S)\epsilon_{0}[\phi_{2}(\mathbf{Z,C})-E\{\phi_{2}(\mathbf{Z,C})|\mathbf{C},S=0\}]:\textit{$\phi_{2}(\mathbf{Z,C})$ unrestricted}\right\}$
@@
where the final equality follows since $f(\mathbf{A}|,\mathbf{Z},S=1,\mathbf{C})=f(\mathbf{A}|\mathbf{Z},S=1,\mathbf{C};\pi^{\dagger})$.
@@
since $f(\mathbf{A,Z}|S=1,\mathbf{C})=f(\mathbf{A,Z}|S=1,\mathbf{C};\pi^{\dagger},\tau^{\dagger})$.
@@
$\Pi_{\mathcal{H}}\left(\mu(\mathbf{Z},S,\mathbf{C})|\{e_{0}(\mathbf{C})+e_{1}(S,\mathbf{C})+e_{2}(\mathbf{Z,C})\}^{\perp}:e_{0}\in\mathcal{C}_{0},e_{1}\in\mathcal{C}_{1},e_{2}\in\mathcal{C}_{2}\right)\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\epsilon^{*}(\psi^{\dagger})$
@@
$\tilde{S}=S-P_{\sigma}(\mathbf{Z,C}).$
@@
×$\times\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\epsilon^{*}$
@@
×$\times\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\epsilon^{*}$
@@
To further simply the previous expression for the efficient score, note firstly that
$P_{\sigma}(\mathbf{Z,C})$ $=\frac{f(S=1|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},1,\mathbf{C})}{f(S=1|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},1,\mathbf{C})+f(S=0|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},0,\mathbf{C})}$
$=\frac{\frac{f(S=1|\mathbf{Z,C})}{f(S=0|\mathbf{Z,C})}\frac{\sigma^{-2}(\mathbf{Z},1,\mathbf{C})}{\sigma^{-2}(\mathbf{Z},0,\mathbf{C})}}{\frac{f(S=1|\mathbf{Z,C})}{f(S=0|\mathbf{Z,C})}\frac{\sigma^{-2}(\mathbf{Z},1,\mathbf{C})}{\sigma^{-2}(\mathbf{Z},0,\mathbf{C})}+1}$
and $1-P_{\sigma}(\mathbf{Z,C})=E\{\sigma^{-2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}^{-1}$. Secondly,
$\frac{E\{\mu(\mathbf{Z},S,\mathbf{C})\sigma^{-2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}}{E\{\sigma^{-2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}}$ $=\frac{\mu(\mathbf{Z},1,\mathbf{C})\frac{f(S=1|\mathbf{Z,C})}{f(S=0|\mathbf{Z,C})}\frac{\sigma^{-2}(\mathbf{Z},1,\mathbf{C})}{\sigma^{-2}(\mathbf{Z},0,\mathbf{C})}+\mu(\mathbf{Z},0,\mathbf{C})}{\frac{f(S=1|\mathbf{Z,C})}{f(S=0|\mathbf{Z,C})}\frac{\sigma^{-2}(\mathbf{Z},1,\mathbf{C})}{\sigma^{-2}(\mathbf{Z},0,\mathbf{C})}+1}$
$=\mu(\mathbf{Z},1,\mathbf{C})P_{\sigma}(\mathbf{Z,C})+\mu(\mathbf{Z},0,\mathbf{C})\{1-P_{\sigma}(\mathbf{Z,C})\}$
$=\delta\mu(\mathbf{Z},\mathbf{C})P_{\sigma}(\mathbf{Z,C})+\mu(\mathbf{Z},0,\mathbf{C})$
@@
$\mu(\mathbf{Z},S,\mathbf{C})-\frac{E\{\mu(\mathbf{Z},S,\mathbf{C})\sigma^{-2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}}{E\{\sigma^{-2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}}=\delta\mu(\mathbf{Z,C})\tilde{S}$
Thirdly,
$\frac{E\{\mu(\mathbf{Z},S,\mathbf{C})\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\tilde{S}|\mathbf{C}\}}{E\{\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\tilde{S}^{2}|\mathbf{C}\}}$
$=\frac{E\left[\mu(\mathbf{Z},1,\mathbf{C})\frac{f(S=1|\mathbf{Z,C})}{f(S=0|\mathbf{Z,C})}\frac{\sigma^{-2}(\mathbf{Z},1,\mathbf{C})}{\sigma^{-2}(\mathbf{Z},0,\mathbf{C})}\{1-P_{\sigma}(\mathbf{Z,C})\}-\mu(\mathbf{Z},0,\mathbf{C})P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right]}{E\left[\frac{f(S=1|\mathbf{Z,C})}{f(S=0|\mathbf{Z,C})}\frac{\sigma^{-2}(\mathbf{Z},1,\mathbf{C})}{\sigma^{-2}(\mathbf{Z},0,\mathbf{C})}\{1-P_{\sigma}(\mathbf{Z,C})\}^{2}+P_{\sigma}(\mathbf{Z,C})^{2}|\mathbf{C}\right]}$
$=\frac{E\left\{\delta\mu(\mathbf{Z},\mathbf{C})P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}{E\left\{P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}$
@@
×$\times\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\epsilon^{*}$
$=\left[\delta\mu(\mathbf{Z,C})-\frac{E\left\{\delta\mu(\mathbf{Z},\mathbf{C})P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}{E\left\{P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}\right]\tilde{S}\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\epsilon^{*}$
$=\left[\delta\mu(\mathbf{Z,C})-\frac{E\left\{\delta\mu(\mathbf{Z},\mathbf{C})P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}{E\left\{P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}\right]$
×$\quad\times P_{\sigma}(\mathbf{Z,C})\{1-P_{\sigma}(\mathbf{Z,C})\}E\{\sigma^{2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}\frac{(-1)^{1-S}}{f(S|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},S,\mathbf{C})}\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\epsilon^{*}$
$=\left[\delta\mu(\mathbf{Z,C})-\frac{E\left\{\delta\mu(\mathbf{Z},\mathbf{C})P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}{E\left\{P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}\right]P_{\sigma}(\mathbf{Z,C})\frac{(-1)^{1-S}f(\mathbf{Z}|\mathbf{C})}{f(S,\mathbf{Z}|\mathbf{C})}\epsilon^{*}.$
@@
$r_{0}^{opt}(\mathbf{Z},S,\mathbf{C})=\left\{\delta\mu(\mathbf{Z,C})-\frac{E\left\{\delta\mu(\mathbf{Z},\mathbf{C})P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}{E\left\{P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}\right\}2(-1)^{1-S}P_{\sigma}(\mathbf{Z,C}).$
@@
confirming that the optimal index function is in the right subspace. Efficiency at the intersection submodel follows from general results in Robins and Rotnitzky, (2001).
@@
$m_{opt}(\mathbf{C})$ $=E\left[\left\{\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\right\}^{2}\sigma^{2}(\mathbf{Z},S,\mathbf{C})\bigg{|}C\right]^{-1}E\left\{U_{\psi^{\dagger}}\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\epsilon^{*}\bigg{|}C\right\}$
$=E\left[\left\{\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\right\}^{2}\sigma^{2}(Z,S,\mathbf{C})\bigg{|}C\right]^{-1}E\left\{E(U_{\psi^{\dagger}}\epsilon^{*}|Z,S,\mathbf{C})\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\bigg{|}C\right\}$
$=E\left[\left\{\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\right\}^{2}\sigma^{2}(Z,S,\mathbf{C})\bigg{|}C\right]^{-1}E\left\{\mu(\mathbf{Z},S,\mathbf{C})\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\bigg{|}C\right\}$
$=E\left[\left\{\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\right\}^{2}\sigma^{2}(Z,S,\mathbf{C})\bigg{|}C\right]^{-1}\{\mu(Z=1,S=1,\mathbf{C})-\mu(Z=1,S=0,\mathbf{C})\}.$
@@
for $U(\mathbf{O};r)=\partial\log\{f(\mathbf{O};r)\}/\partial r$, where $f(\mathbf{O};r)$ is a one-dimensional regular parametric submodel of $\mathcal{M}_{np}$. By the product rule,

### v6

@@ A.10 The average treatment effect in the treated
$\beta(\mathbf{C})$ $\equiv\frac{\delta^{Y}(\mathbf{C})-t(1,\mathbf{C})}{\delta^{A}(\mathbf{C})}$
$\delta^{Y}(\mathbf{C})$ $\equiv E(Y|Z=1,S=1,\mathbf{C})-E(Y|Z=0,S=1,\mathbf{C})$
$\delta^{A}(\mathbf{C})$ $\equiv f(A=1|Z=1,S=1,\mathbf{C})-f(A=1|Z=0,S=1,\mathbf{C}).$
@@ Appendix B Proofs
$E(Y|A=a,z,S=1,\mathbf{c})=$ $\beta(a,z,\mathbf{c})+\gamma(z,\mathbf{c})\{a-f(A=1|Z=z,S=1,\mathbf{c})\}$
@@
$t(1,\mathbf{c})$ $=E(Y^{0}|Z=1,S=0,\mathbf{C}=\mathbf{c})-E(Y^{0}|Z=0,S=0,\mathbf{C}=\mathbf{c})$ (Assumption 4)
@@
$=\beta(a,z,\mathbf{c})+\gamma(z,\mathbf{c})\{a-f(A=1|Z=z,S=1,\mathbf{c})\}$
@@
$\epsilon_{1}$ $\equiv Y-\beta(\mathbf{A,Z,C})-[q(\mathbf{A,Z,C})-E\{q(\mathbf{A,Z,C})|\mathbf{Z},S=1,\mathbf{C}\}]-b^{*}_{1}(\mathbf{C})-t(\mathbf{Z,C})$
$\epsilon^{*}_{1}$ $\equiv\epsilon_{1}+[q(\mathbf{A,Z,C})-E\{q(\mathbf{A,Z,C})|\mathbf{Z},S=1,\mathbf{C}\}]$
$\epsilon_{0}$ $\equiv Y-b_{0}(\mathbf{C})-t(\mathbf{Z,C})$
@@
$c_{2}(\mathbf{A,Z},\mathbf{C})=[\phi_{1}(\mathbf{Z,C})-E\{\phi_{1}(\mathbf{Z,C})|S=1,\mathbf{C}\}][q(\mathbf{A,Z,C})-E\{q(\mathbf{A,Z,C})|\mathbf{Z},S=1,\mathbf{C}\}].$
@@
=$=\bigg{\{}S\epsilon^{*}_{1}[\phi_{1}(\mathbf{Z,C})-E\{\phi_{1}(\mathbf{Z,C})|S=1,\mathbf{C}\}]:\textrm{$\phi_{1}(\mathbf{Z,C})$ unrestricted}\bigg{\}}\cap L_{2}^{0}.$
@@
$\Lambda^{\perp}_{nuis_{2}}\cap\Lambda^{\perp}_{nuis_{8}}=\left\{(1-S)\epsilon_{0}[\phi_{2}(\mathbf{Z,C})-E\{\phi_{2}(\mathbf{Z,C})|\mathbf{C},S=0\}];\textrm{$\phi_{2}(\mathbf{Z,C})$ unrestricted}\right\}.$
@@
$\quad+(1-S)\epsilon_{0}[\phi_{2}(\mathbf{Z,C})-E\{\phi_{2}(\mathbf{Z,C})|\mathbf{C},S=0\}]:\textrm{$\phi_{1}(\mathbf{Z,C}),\phi_{2}(\mathbf{Z,C})$ unrestricted}\bigg{\}}.$
@@
$=E\left[\phi(\mathbf{Z},S,\mathbf{C};\tau^{\dagger},\alpha^{\dagger},\rho^{\dagger})\{b_{0}(\mathbf{C})+b_{1}(\mathbf{C})S+t(\mathbf{Z,C})-b_{0}(\mathbf{C};\theta_{0}^{\dagger})-b_{1}(\mathbf{C};\theta_{1}^{\dagger})S-t(\mathbf{Z,C};\nu^{\dagger})\}\right]=0.$
@@
where the final equality follows since $f(\mathbf{Z}|S,\mathbf{C})=f(\mathbf{Z}|S,\mathbf{C};\tau^{\dagger},\rho^{\dagger})$.
@@
where the final equality follows since $f(S|\mathbf{Z,C})=f(S|\mathbf{Z,C};\alpha^{\dagger},\rho^{\dagger})$.
@@
$E(Y|A=a,Z=z,S=1,\mathbf{c})$
$=E(Y^{1}-Y^{0}|A=1,Z=z,S=1,\mathbf{c})a$
$\quad+\gamma(z,\mathbf{c})\{a-Pr(A=1|Z=z,S=1,\mathbf{c})\}$
@@
$E\{Y-E(Y|Z=z,S=0,\mathbf{c})+E(Y|Z=0,S=0,\mathbf{c})|Z=z,S=1,\mathbf{c}\}$
@@
$\Lambda^{\perp}_{nuis_{2}}\cap\Lambda^{\perp}_{nuis_{8}}=\left\{(1-S)\epsilon_{0}[\phi_{2}(\mathbf{Z,C})-E\{\phi_{2}(\mathbf{Z,C})|\mathbf{C},S=0\}]:\textrm{$\phi_{2}(\mathbf{Z,C})$ unrestricted}\right\}$
@@
where the final equality follows since $f(\mathbf{A}|\mathbf{Z},S=1,\mathbf{C})=f(\mathbf{A}|\mathbf{Z},S=1,\mathbf{C};\pi^{\dagger})$.
@@
since $f(\mathbf{A,Z}|S=1,\mathbf{C})=f(\mathbf{A,Z}|S=1,\mathbf{C};\pi^{\dagger},\tau^{\dagger},\rho^{\dagger})$.
@@
$=\Pi_{\mathcal{H}}\left(\mu(\mathbf{Z},S,\mathbf{C})|\{e_{0}(\mathbf{C})+e_{1}(S,\mathbf{C})+e_{2}(\mathbf{Z,C})\}^{\perp}:e_{0}\in\mathcal{C}_{0},e_{1}\in\mathcal{C}_{1},e_{2}\in\mathcal{C}_{2}\right)\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\epsilon^{*}(\psi^{\dagger})$
@@
$\tilde{S}$ $\equiv S-\frac{E\{S\sigma^{-2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}}{E\{\sigma^{-2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}}$
$=S-\frac{f(S=1|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},1,\mathbf{C})}{f(S=1|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},1,\mathbf{C})+f(S=0|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},0,\mathbf{C})}.$
@@
×$\times\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\epsilon^{*}(\psi^{\dagger})$
@@
×$\times\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\epsilon^{*}(\psi^{\dagger})$
@@
To further simply the previous expression for the efficient score, note that
$\mu(\mathbf{Z},S,\mathbf{C})$ $=\mu(\mathbf{Z},1,\mathbf{C})S,$
$\frac{E\{\mu(\mathbf{Z},S,\mathbf{C})\sigma^{-2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}}{E\{\sigma^{-2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}}$ $=\frac{\mu(\mathbf{Z},1,\mathbf{C})f(S=1|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},1,\mathbf{C})}{f(S=1|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},1,\mathbf{C})+f(S=0|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},0,\mathbf{C})}$
@@
$\mu(\mathbf{Z},S,\mathbf{C})-\frac{E\{\mu(\mathbf{Z},S,\mathbf{C})\sigma^{-2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}}{E\{\sigma^{-2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}}=\mu(\mathbf{Z},1,\mathbf{C})\tilde{S}.$
Further,
$\frac{E\{\mu(\mathbf{Z},S,\mathbf{C})\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\tilde{S}|\mathbf{C}\}}{E\{\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\tilde{S}^{2}|\mathbf{C}\}}=\frac{E\left\{\mu(\mathbf{Z},1,\mathbf{C})P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}{E\left\{P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}.$
@@
×$\times\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\epsilon^{*}(\psi^{\dagger})$
$=\left[\mu(\mathbf{Z},1,\mathbf{C})-\frac{E\left\{\mu(\mathbf{Z},1,\mathbf{C})P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}{E\left\{P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}\right]\tilde{S}\sigma^{-2}(\mathbf{Z},S,\mathbf{C})\epsilon^{*}(\psi^{\dagger})$
$=\left[\mu(\mathbf{Z},1,\mathbf{C})-\frac{E\left\{\mu(\mathbf{Z},1,\mathbf{C})P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}{E\left\{P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}\right]P_{\sigma}(\mathbf{Z,C})\frac{(-1)^{1-S}}{f(S|\mathbf{Z},\mathbf{C})}\epsilon^{*}(\psi^{\dagger})$
$=\left[\mu(\mathbf{Z},1,\mathbf{C})-\frac{E\left\{\mu(\mathbf{Z},1,\mathbf{C})P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}{E\left\{P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}\right]P_{\sigma}(\mathbf{Z,C})\frac{(-1)^{1-S}f(\mathbf{Z}|\mathbf{C})}{f(S,\mathbf{Z}|\mathbf{C})}\epsilon^{*}(\psi^{\dagger}).$
@@
$r_{0}^{opt}(\mathbf{Z},S,\mathbf{C})=\left\{\mu(\mathbf{Z},1,\mathbf{C})-\frac{E\left\{\mu(\mathbf{Z},1,\mathbf{C})P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}{E\left\{P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}\right\}2(-1)^{1-S}P_{\sigma}(\mathbf{Z,C}).$
@@
confirming that the optimal index function is in the correct subspace. Efficiency at the intersection submodel follows from general results in Robins and Rotnitzky, (2001).
@@
$m_{opt}(\mathbf{C})$ $=E\left[\left\{\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\right\}^{2}\sigma^{2}(Z,S,\mathbf{C})\bigg{|}C\right]^{-1}E\left\{U_{\psi^{\dagger}}\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\epsilon^{*}(\psi^{\dagger})\bigg{|}C\right\}$
$=E\left[\left\{\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\right\}^{2}\sigma^{2}(Z,S,\mathbf{C})\bigg{|}C\right]^{-1}E\left\{E(U_{\psi^{\dagger}}\epsilon^{*}(\psi^{\dagger})|Z,S,\mathbf{C})\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\bigg{|}C\right\}$
$=E\left[\left\{\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\right\}^{2}\sigma^{2}(Z,S,\mathbf{C})\bigg{|}C\right]^{-1}E\left\{\mu(Z,S,\mathbf{C})\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\bigg{|}C\right\}$
$=E\left[\left\{\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\right\}^{2}\sigma^{2}(Z,S,\mathbf{C})\bigg{|}C\right]^{-1}\{\mu(Z=1,S=1,\mathbf{C})-\mu(Z=0,S=1,\mathbf{C})\}.$
@@
for $U(\mathbf{O};r)=\partial\log\{f(\mathbf{O};r)\}/\partial r$, where $f(\mathbf{O};r)$ is a one-dimensional regular parametric submodel of the nonparametric model for the observed data distribution. By the product rule,